OpenStudy (anonymous):

Need help with factoring quadratics, will medal. Factor completely 81x2– 100

OpenStudy (mayankdevnani):

\[\large \bf 81x^2-100\] you can write it as also :- \[\large \bf 81x^2-10^2\]

OpenStudy (mayankdevnani):

Use \[\large \bf a^2-b^2=(a-b)(a+b)\]

OpenStudy (mayankdevnani):

and then solve

OpenStudy (mayankdevnani):

hope you understand. @AskingAlex_cx

OpenStudy (anonymous):

So plug in for a^2 - b^2... 81x^2-10^2 right?

OpenStudy (mayankdevnani):

right

OpenStudy (mayankdevnani):

can you solve it ? @AskingAlex_cx

OpenStudy (anonymous):

Help me understand @mayankdevnani

OpenStudy (mayankdevnani):

ok

OpenStudy (mayankdevnani):

\[\large \bf 81x^2-10^2\] Use \[\large \bf a^2-b^2=(a-b)(a+b)\] we get, \[\large \bf (9x)^2-(10)^2\] \[\large \bf (9x-10)(9x+10)\]

OpenStudy (mayankdevnani):

there you go ! your problem is SOLVED !

OpenStudy (anonymous):

I just didn't know where to take it next, I thought about sqrt the 81, but wouldm't you have to do the same to both sides?

OpenStudy (mayankdevnani):

see 81=9*9=9^2

OpenStudy (mayankdevnani):

and 100=10*10=10^2

OpenStudy (anonymous):

yes, but the 81 is already squared, so square rooting would make it just 9 right?

OpenStudy (mayankdevnani):

no 81 is not squared,x is squared. Notice in your question DEEPLY. `You have put squared on x not on 81`

OpenStudy (anonymous):

ohh thanks a lot!! @mayankdevnani

OpenStudy (mayankdevnani):

welcome :)